OpenStudy (anonymous):

Solve for x: i got lost in the math 2x/(x -4)+(5/(x + 4))=(2x^2 + 1)/(x^2 - 16)

OpenStudy (callisto):

\[\frac{2x}{(x +4)}+\frac{5}{(x + 4)}=\frac{(2x^2 + 1)}{(x^2 - 16)}\]\[\frac{2x+5}{(x +4)}=\frac{(2x^2 + 1)}{(x+4)(x-4)}\]\[2x+5=\frac{(2x^2 + 1)}{(x-4)}\]\[(2x+5)(x-4)=(2x^2 + 1)\]\[(2x^2+5x-8x-20)=(2x^2 + 1)\]\[-3x=21\]\[x=-7\]

OpenStudy (anonymous):

im sorry its supposed to be 2/(x-4) :\

OpenStudy (callisto):

Oh... my bad.... I read something wrong... Gimme some time.. I'll re-type that again :S

OpenStudy (anonymous):

thank you very much

OpenStudy (callisto):

\[\frac{2x}{x-4} + \frac{5x}{x+4} = \frac{2x^2 + 1}{x^2 - 16}\]\[\frac{2x(x+4)}{(x-4)(x+4)} + \frac{5x(x-4)}{(x+4)(x-4)} = \frac{2x^2 + 1}{x^2 - 16}\]\[\frac{2x(x+4)}{(x^2-16)} + \frac{5x(x-4)}{(x^2-16)} = \frac{2x^2 + 1}{x^2 - 16}\]\[2x(x+4) + 5x(x-4)= 2x^2 + 1\]\[2x^2+8x + 5x^2 -20x= 2x^2 + 1\]\[5x^2 -12x-1=0\] I think you'd get something ugly :|

OpenStudy (anonymous):

i agree, lol, im just going to guess, thank you for your help :)

OpenStudy (callisto):

Welcome. Use quadratic formula to solve